# Failure of famous functional inequalities on Finsler manifolds: the influence of $S$-curvature

### v1

@@ 5.1. Infinite reversibility
In this subsection we prove Theorem 1.2, which establish a surprising relationship between the $S$-curvature and reversibility. To do this, we recall the following Hardy inequality.
@@
(Zhao [36, Theorem 3.8]) Let $(M,F,\operatorname{\mathfrak{m}})$ be an $n$-dimensional forward complete noncompact $\operatorname{\mathsf{FMMM}}$ with
@@
[…]
@@
[…]
@@ 6. Examples & Counterexamples: Funk metric spaces
The purpose of this section is twofold. First, in subsection 6.1 we investigate Funk metric spaces, a large class of non-Riemannian Finsler manifolds that provide the geometric background for the applicability of our main results, Theorems 1.1-1.2. Second, in subsection 6.2, as we already noticed in the Introduction, we show that the combined assumption on the Ricci and $S$-curvatures cannot be replaced by the weighted Ricci curvature ${\bf Ric}_{N}$, in spite of the fact that the latter notion is a combination of $\mathbf{Ric}$ and $\mathbf{S}$.
@@
6.1. Funk metric spaces: applicability of Theorems 1.1-1.2
@@ A Finsler metric $F$ on $\Omega$ is called a (general) Funk metric if
Any $n$-dimensional Funk metric space $(\Omega,F,\operatorname{\mathfrak{m}}_{BH})$ endowed with the Busemann–Hausdorff measure verifies the assumptions of Theorems 1.1-1.2. In particular, the Hardy inequality, Heisenberg–Pauli–Weyl uncertainty principle and Caffarelli–Kohn–Nirenberg inequality fail on $(\Omega,F,\operatorname{\mathfrak{m}}_{BH})$, respectively.
Before giving the proof of Theorem 6.1, we recall/collect some useful properties of Funk metric spaces. We first notice that every Funk metric is projectively flat. In addition, according to Li [19] and Shen [31], a Finsler metric $F$ on $\Omega$ is a Funk metric if and only if there is a unique Minkowski norm $\phi=\phi(y)$ on $\mathbb{R}^{n}$ such that
@@
Let $(\Omega,F)$ be a Funk metric space and let $\phi=\phi(y)$ the Minkowski norm satisfying (6.1). Then the following statements hold::::
@@
    $(\Omega,F)$ is forward complete but not backward complete with $\lambda_{F}(\Omega)=+\infty$. In particular,
@@
$\partial\Omega\ni z:=x+\frac{y}{F(x,y)}=c(s)+\frac{y}{F(c(s),\dot{c}(s))}=x+sy+\frac{y}{F(c(s),\dot{c}(s))},$
@@
$F(c(s),\dot{c}(s))=\frac{F(x,y)}{1-F(x,y)s}=\frac{F(y)}{1-F(y)s},$
where $F(y)$ is $F(x,y)$ for short. Thus, formula (6.3) giving the distance ${d}_{F}(x_{1},x_{2})$ between $x_{1}=c(s_{1})$ to $x_{2}=c(s_{2})$ is
${d}_{F}(x_{1},x_{2})=$ $\int_{s_{1}}^{s_{2}}F(c(s),\dot{c}(s)){\operatorname{d}}s=\int_{s_{1}}^{s_{2}}\frac{F(y)}{1-F(y)s}{\operatorname{d}}s=\ln\left[\frac{1-F(y)s_{1}}{1-F(y)s_{2}}\right]=\ln\left[\frac{\frac{1}{F(y)}-s_{1}}{\frac{1}{F(y)}-s_{2}}\right]$
@@
On the other hand, when $x_{1}=x$ and $x_{2}=\mathbf{0}$, from (6.6) we have $\lambda=\frac{1}{\phi(-x)}+1$; thus (6.7) gives
@@
Let $t\mapsto c(t)$, $t\in[0,1],$ be the unit speed geodesic from $c(0)=x_{1}$ to $c(1)=x_{2}$ (i.e., the ray $l_{x_{1}x_{2}}$). We extend $c$ to the point $z=x_{1}+\lambda(x_{2}-x_{1})\in\partial\Omega$; hence,
@@
$d_{F}(c(0),c(t))=d_{F}(x_{1},c(t))\geq-\ln(1-\phi(c(t)))-\ln(1+\phi(x_{1}))\rightarrow+\infty,$
as $\phi(c(t))\rightarrow 1^{-}$, which means that every unit speed geodesic can be forward extended to $t=+\infty$. Thus, $(\Omega,F)$ is forward complete. On the other hand, (6.8) implies $d_{F}(x,\mathbf{0})\rightarrow\ln 2$ as $\phi(-x)\rightarrow 1^{-}$. Hence, ($\Omega$, $F$) cannot be backward complete because the bounded closed backward ball $\overline{B^{-}_{\mathbf{0}}(\ln 2)}=\Omega$ is not compact.
@@
with equality if and only if $\xi=\alpha\mathfrak{L}_{\phi}(x)$ for some $\alpha\geq 0$; here, $\mathfrak{L}_{\phi}$ stands for the Legendre transformation induced by $\phi$. It costs no generality to assume that $\phi^{*}(\mathfrak{L}_{\phi}(x))\leq\phi^{*}(-\mathfrak{L}_{\phi}(x))$. According to [11, 28], the co-metric $F^{*}$ of $F$ is given by
@@
where $\phi^{*}$ is the co-metric of $\phi$. Therefore, by (2.5) one has
@@
In particular, $\lambda_{F}(x)\to+\infty$ as $\phi(x)\rightarrow 1^{-}$, thus $\lambda_{F}(\Omega)=+\infty$. Moreover,
@@
By Proposition 6.1/(iii), $(\Omega,F)$ is forward complete. In addition, according to Shen [26], one has that $(\Omega,F)$ is a Cartan–Hadamard manifold with
@@
In particular, we may choose
@@
in the assumptions of Theorems 1.1 &\&& 1.2, which guarantee the validity of (1.5) and (1.7). ∎
@@
We conclude this subsection by recalling the interpolation metric considered in Kristály and Rudas [18]; namely, for $a\in[0,1]$, let $F_{a}:T\mathbb{B}^{n}\to[0,\infty)$ be given by
@@
It is clear that $F_{a}$ interpolates between the Funk metric (1.6) for $a=1$ with ${\bf K}=-1/4$ and ${\bf S}_{a}=(n+1)/2,$ and the Hilbert/Klein metric for $a=0$ with ${\bf K}=-1$ and ${\bf S}_{a}=0$, where ${\bf S}_{a}$ denotes the $S$-curvature of the Busemann–Hausdorff measure $\operatorname{\mathfrak{m}}_{a}$. Note that $(\mathbb{B}^{n},F_{a})$ is a Randers-type projectively flat Finsler space with the reversibility $\lambda_{F_{a}}(\mathbb{B}^{n})=\frac{1+a}{1-a}$ for $a\in[0,1)$ and $\lambda_{F_{1}}(\mathbb{B}^{n})=+\infty$. Moreover, in Kajántó and Kristály [14, Theorems 1.1&1.2] – up to an isometry – it is proved that the flag curvature and $S$-curvature of $F_{a}$ with $a\in(0,1)$, can be estimated as
@@
moreover, these estimates are sharp for every $a\in(0,1)$. In particular, for every $a\in[0,1)$ one has that $\inf_{(x,y)\in T\mathbb{B}^{n}}{\bf S}_{a}(x,y)=0$, thus Theorem 1.2 does not apply (i.e., $0=h<k$). However, by using Theorem 6.1, elementary estimates as in [18] and the validity of famous functional inequalities on Riemannian Cartan–Hadamard manifolds, see [10, 16], we can state:
@@
Let $a\in[0,1]$, and consider the interpolated Funk-type metric space $(\mathbb{B}^{n},F_{a},\operatorname{\mathfrak{m}}_{a})$ endowed with its natural Busemann–Hausdorff measure $\operatorname{\mathfrak{m}}_{a}$. Then the Hardy inequality, Heisenberg–Pauli–Weyl uncertainty principle and Caffarelli–Kohn–Nirenberg inequality hold on $(\mathbb{B}^{n},F_{a},\operatorname{\mathfrak{m}}_{a})$ – with possibly not sharp constants – if and only if $a\in[0,1)$.

### v2

@@ 5.1. Infinite reversibility
In this subsection we prove Theorem 1.2, which establish a surprising relationship between the $S$-curvature and reversibility. To do this, we recall the following Hardy inequality established by Zhao [36, Theorem 3.8].
@@
(Zhao [36]) Let $(M,F,\operatorname{\mathfrak{m}})$ be an $n$-dimensional forward complete noncompact $\operatorname{\mathsf{FMMM}}$ with
@@
Given a fixed point $o\in M$, set ${\mathop{r}\limits^{\leftarrow}}(x):=d_{F}(x,o)$. Then for any $p\in(1,n)$ and $\beta\in(-n,\infty)$,
$\int_{M}\max\{F^{*p}(\pm{\operatorname{d}}u)\}{\mathop{r}\limits^{\leftarrow}}^{\beta+p}\operatorname{d\mathfrak{m}}\geq\left(\frac{n+\beta}{p}\right)^{p}\int_{M}|u|^{p}{\mathop{r}\limits^{\leftarrow}}^{\beta}\operatorname{d\mathfrak{m}},\quad\forall\,u\in C^{\infty}_{0}(M).$
@@
Suppose by contradiction that $\lambda_{F}(M)$ is finite. Choosing an arbitrary point $o\in M$, set $r(x)=d_{F}(o,x)$ and ${\mathop{r}\limits^{\leftarrow}}(x):=d_{F}(x,o)$. Then (2.8) and (2.5) yield
${\lambda^{-1}_{F}(M)}\,r\leq{\mathop{r}\limits^{\leftarrow}}\leq\lambda_{F}(M)\,r,\quad\max\{F^{*}(\pm{\operatorname{d}}u)\}\leq\lambda_{F}(M)\,F^{*}({\operatorname{d}}u).$ (5.1)
@@
5.2. Topological properties of projectively flat manifolds
In this subsection we prove Theorem 1.3, which provides another surprising phenomenon coming from the behavior of the $S$-curvature. Throughout this subsection, let $\Omega$ be a non-empty domain (i.e., connected open subset) in $\mathbb{R}^{n}$ with $\mathbf{0}\in\Omega$. A Finsler metric $F$ on $\Omega$ is said to be projectively flat if the images of all geodesics are straight lines. Analytically, this property can be characterized as follows.
Lemma 5.1.
(Shen [27]) The metric $F$ on $\Omega$ is projectively flat if and only if the geodesic coefficients defined by (2.7) satisfy $G^{i}=Py^{i}$, where
$P:=\frac{y^{k}}{2F}\frac{\partial F}{\partial{x^{k}}}$ (5.2)
is called the projective factor of $F$.
Remark 5.1.
The projective factor $P$ is positively $1$-homogenous, i.e., $P(x,\alpha y)=\alpha P(x,y)$ for any $\alpha>0$. Then Euler’s theorem (cf. [2, Theorem 1.2.1]) yields
$y^{m}\frac{\partial P}{\partial y^{m}}(x,y)=P(x,y).$ (5.3)
In the following result we prove that the completeness of a projectively flat Finsler manifold is deeply influenced by the $S$-curvature.
Theorem 5.2.
Let $(\Omega,F,\mathscr{L}^{n})$ be an $n$-dimensional projectively flat Finsler manifold endowed with the Lebesgue measure. Assume that $\Omega$ is bounded in $\mathbb{R}^{n}$. Then $(\Omega,F)$ cannot be backward ((resp., forward)) complete whenever $\mathbf{S}\geq 0$ ((resp., $\mathbf{S}\leq 0$).).
Proof.
We only consider $\mathbf{S}\geq 0$; the other case can be treated similarly. Let $(x^{i})$ be the standard coordinate system of $\mathbb{R}^{n}$. It follows by (2.11), Lemma 5.1 and (5.3) that
$0\leq\mathbf{S}(x,y)=\frac{\partial G^{m}(x,y)}{\partial y^{m}}=\frac{\partial(P(x,y)y^{m})}{\partial y^{m}}=(n+1){P(x,y)},$
which implies
$P(x,y)\geq 0,\quad\forall\,(x,y)\in T\Omega\backslash\{0\}.$ (5.4)
As $\mathbf{0}\in\Omega$ and $F$ is projectively flat, a geodesic $t\mapsto\gamma(t)$ starting from a point $\mathbf{x}_{o}\in\Omega$ in the direction $\mathbf{a}\in\partial\Omega$ has the form
$\gamma(t)=\mathbf{x}_{o}+f(t)\mathbf{a},\quad t\in(-\varepsilon,\varepsilon),$
with $f(0)=0$ and $f^{\prime}(t)>0$. The existence of $\varepsilon>0$ follows from the standard theory of ODE, as $\gamma$ satisfies the geodesic equation (2.6).
Assume by contradiction that $(\Omega,F,\operatorname{\mathfrak{m}})$ is backward complete; then $\gamma$ can be extended to $t\in(-\infty,\varepsilon)$, which implies $f\in C^{2}((-\infty,\varepsilon))$ (in fact, $f$ is smooth). Since $\gamma$ satisfies (2.6), it follows by [2, Exercise 5.2.4/(d)] that $\gamma$ has constant velocity, i.e.,
$F(\gamma(t),\gamma^{\prime}(t))=F(\gamma(t),f^{\prime}(t)\mathbf{a})=\text{const}=:c,$ (5.5)
which means $f^{\prime}(t)>0$ for all $t\in(-\infty,\varepsilon)$. Due to Lemma 5.1, the geodesic equation (2.6) reduces to
$f^{\prime\prime}(t)+2P\left(\gamma(t),f^{\prime}(t)\mathbf{a}\right)f^{\prime}(t)=0,$ (5.6)
which together with (5.4), Remark 5.1 and $f^{\prime}(t)>0$ yields
$f^{\prime\prime}(t)=-2\left(f^{\prime}(t)\right)^{2}P(\gamma(t),\mathbf{a})\leq 0.$
Therefore, by a Taylor expansion, it follows that $f$ satisfies
$f(t)\leq f(0)+f^{\prime}(0)(t-0)=f^{\prime}(0)t,\quad\forall\,t\in(-\infty,\varepsilon).$ (5.7)
Note that the image of $\gamma(t)=\mathbf{x}_{o}+f(t)\mathbf{a}$ is the straight line $l(s)=\mathbf{x}_{o}+s\mathbf{a}$. Since $\Omega$ is bounded, we may assume that $l(s)=\mathbf{x}_{o}+s\mathbf{a}$ intersects $\partial\Omega$ in the opposite direction of $\mathbf{a}$ at some $s_{-}\in(-\infty,0)$. Then the Euclidean distance satisfies
$|f(t)\mathbf{a}|=|\gamma(t)-\mathbf{x}_{o}|\leq|l(s_{-})-\mathbf{x}_{o}|=|s_{-}\mathbf{a}|,\quad\forall\,t\in(-\infty,0),$
which combined with (5.7) implies
$\infty<s_{-}\leq f(t)\leq f^{\prime}(0)t,\quad\forall\,t\in(-\infty,0).$
Since $f^{\prime}(0)>0$, the latter estimate provides a contradiction. Then $(\Omega,F,\operatorname{\mathfrak{m}})$ is not backward complete. ∎
Theorem 5.2 is optimal in the sense that the boundedness of $\Omega$ in $\mathbb{R}^{n}$ cannot be canceled. Indeed, the following two examples point out that in both cases (i.e., $\mathbf{S}=0$ and $\mathbf{S}>0$), if $\Omega$ is unbounded in $\mathbb{R}^{n}$, the corresponding projectively flat manifold can be complete, i.e., both forward and backward complete.
Example 5.1.
Let $(\Omega,F,\mathscr{L}^{n}):=(\mathbb{R}^{n},|\cdot|,\mathscr{L}^{n})$ be the Euclidean space endowed with the Lebesgue measure. Obviously, $\Omega$ is unbounded in $\mathbb{R}^{n}$; although $(\Omega,F,\mathscr{L}^{n})$ is projectively flat with $\mathbf{S}=0$, it is complete.
Example 5.2.
Let $(\Omega,F,\mathscr{L}^{n}):=(\mathbb{R}^{n},F,\mathscr{L}^{n})$, where
$F(x,y):=|y|+\frac{\langle x,y\rangle}{2\sqrt{1+|x|^{2}}}.$ (5.8)
According to Lemma 5.1, a direct computation shows the projective flatness of $F$ and
$P=\frac{1}{2F}\frac{\partial F}{\partial x^{k}}y^{k}=\frac{(1+|x|^{2})|y|^{2}-\langle x,y\rangle^{2}}{4(1+|x|^{2})^{\frac{3}{2}}\Big(|y|+\frac{\langle x,y\rangle}{2\sqrt{1+|x|^{2}}}\Big)}>0,\quad\mathbf{S}={(n+1)P}>0.$
As $F$ is projectively flat, the minimal geodesic from $\mathbf{0}$ to $x$ has the same image of the straight line $\gamma(t)=tx$, $t\in[0,1]$. Then the distances ${d}_{F}(\mathbf{0},x)$ and ${d}_{F}(x,\mathbf{0})$ can be obtained by
${d}_{F}(\mathbf{0},x)$ $=\int^{1}_{0}F(tx,x){\operatorname{d}}t=\int^{1}_{0}|x|+\frac{t|x|^{2}}{2\sqrt{1+t^{2}|x|^{2}}}{\operatorname{d}}t=|x|-\frac{1}{2}+\frac{1}{2}\sqrt{1+|x|^{2}},$
${d}_{F}(x,\mathbf{0})$ $=\int^{1}_{0}F(tx,-x){\operatorname{d}}t=\int^{1}_{0}|x|-\frac{t|x|^{2}}{2\sqrt{1+t^{2}|x|^{2}}}{\operatorname{d}}t=|x|+\frac{1}{2}-\frac{1}{2}\sqrt{1+|x|^{2}}.$
Given $x_{1},x_{2}\in\mathbb{R}^{n}$, the triangle inequality implies
$d_{F}(x_{1},x_{2})\geq d_{F}(\mathbf{0},x_{2})-d_{F}(\mathbf{0},x_{1})=|x_{2}|+\frac{1}{2}\sqrt{1+|x_{2}|^{2}}-|x_{1}|-\frac{1}{2}\sqrt{1+|x_{1}|^{2}},$ (5.9)
$d_{F}(x_{1},x_{2})\geq d_{F}(x_{1},\mathbf{0})-d_{F}(x_{2},\mathbf{0})=|x_{1}|-\frac{1}{2}\sqrt{1+|x_{1}|^{2}}-|x_{2}|+\frac{1}{2}\sqrt{1+|x_{2}|^{2}}.$ (5.10)
Let $c$ be the unit-speed geodesic from $x_{1}$ to $x_{2}$. By the projective flatness, $c$ can be extended beyond $x_{2}$. In fact, $|c(t)|\rightarrow+\infty$ as $t\to\infty,$ since the image of $t\mapsto c(t)$ is a straight line and $\Omega=\mathbb{R}^{n}$. Then (5.9) yields
$d_{F}(c(0),c(t))=d_{F}(x_{1},c(t))\geq|c(t)|+\frac{1}{2}\sqrt{1+|c(t)|^{2}}-|x_{1}|-\frac{1}{2}\sqrt{1+|x_{1}|^{2}}\rightarrow+\infty,$
as $|c(t)|\rightarrow+\infty$, which means that every unit-speed geodesic can be extended forward to $t=+\infty$. Hence, $(\mathbb{R}^{n},F,\mathscr{L}^{n})$ is forward complete.
In a similar manner, $c$ can be also extended before $x_{1}$. By (5.10), we have
$d_{F}(c(t),c(0))=d_{F}(c(t),x_{1})\geq|c(t)|-\frac{1}{2}\sqrt{1+|c(t)|^{2}}-|x_{1}|+\frac{1}{2}\sqrt{1+|x_{1}|^{2}}\rightarrow+\infty,$
as $|c(t)|\rightarrow+\infty$, which means that every unit-speed geodesic can be extended backward to $t=$-$\infty$. Then $(\mathbb{R}^{n},F,\mathscr{L}^{n})$ is backward complete. In conclusion, $(\mathbb{R}^{n},F,\mathscr{L}^{n})$ is complete.
Moreover, a direct computation shows that
$\lambda_{F}(\mathbb{R}^{n})=\sup_{x\in\mathbb{R}^{n}}\sup_{y\in T_{x}\mathbb{R}^{n}\setminus\{0\}}\frac{F(x,-y)}{F(x,y)}=\sup_{x\in\mathbb{R}^{n}}\sup_{y\in\mathbb{R}^{n}\setminus\{0\}}\frac{|y|-\frac{\langle x,y\rangle}{2\sqrt{1+|x|^{2}}}}{|y|+\frac{\langle x,y\rangle}{2\sqrt{1+|x|^{2}}}}=3.$
A direct consequence of Theorem 5.2 reads as follows:
Corollary 5.2.
Let $(\Omega,F,\mathscr{L}^{n})$ be an $n$-dimensional forward complete projectively flat Finsler manifold endowed with the Lebesgue measure. Assume that $\Omega$ is bounded in $\mathbb{R}^{n}$ and $\mathbf{S}\geq 0$. Then $\lambda_{F}(M)=+\infty$.
Having Theorem 5.2, we now have:
Proof of Theorem 1.3.
The proof is a modification of that of Theorem 5.2. Hence, we just give a sketch. Since $(\Omega,F)$ is forward complete, a geodesic $\gamma$ starting from $\mathbf{0}\in\Omega$ in the direction $\mathbf{a}\in\mathbb{S}^{n-1}$ has the form
$\gamma(t)=f(t)\mathbf{a},\quad t\in(-\varepsilon,+\infty),$ (5.11)
with $f(0)=0$ and $f^{\prime}(t)>0$. The assumption $\mathbf{S}\geq(n-1)h>0$ implies
$(n+1)P(x,y)=\mathbf{S}(x,y)\geq(n-1)hF(x,y),\quad\forall\,(x,y)\in T\Omega\backslash\{0\},$
which together with (5.5) and (5.6) yields
$f^{\prime\prime}(t)=-2P\left(\gamma(t),\gamma^{\prime}(t)\right)f^{\prime}(t)\leq-\frac{2(n-1)}{n+1}hcf^{\prime}(t),$
where $c=F(\gamma(t),\gamma^{\prime}(t))$ is the constant speed of $\gamma$. Since $f^{\prime}(t)>0$, the above inequality can be written into
$\ln f^{\prime}(t)]^{\prime}\leq-\frac{2(n-1)}{n+1}hc,$
which yields
$f^{\prime}(t)\leq e^{-\frac{2(n-1)}{n+1}hct}f^{\prime}(0).$
Integrating the above inequality again from 0 to $t$ yields
$f(t)\leq-\frac{(n+1)f^{\prime}(0)}{2(n-1)hc}e^{-\frac{2(n-1)}{n+1}hct}+\frac{(n+1)f^{\prime}(0)}{2(n-1)hc}.$
Then we have
$f(t)\leq\frac{(n+1)f^{\prime}(0)}{2(n-1)hc}=:R<+\infty,\quad\forall\,t\in[0,+\infty).$
This inequality together with (5.11), the forward completeness, and the projective flatness of $(\Omega,F)$ implies
$\Omega\subset\mathbb{B}^{n}(R):=\{y\in\mathbb{R}^{n}\,:\,|y|\leq R\},$
i.e., $\Omega$ is bounded in $\mathbb{R}^{n}$. By Theorem 5.2, $(\Omega,F)$ is not backward complete.
Moreover, we have $\lambda_{F}(\Omega)=+\infty$; indeed, if $\lambda_{F}(\Omega)$ is finite, $(\Omega,F)$ would be also backward complete, coming from its forward completeness. ∎
@@ 6. Examples & Counterexamples: Funk metric spaces
The purpose of this section is twofold. First, in subsection 6.1 we investigate Funk metric spaces, a large class of non-Riemannian Finsler manifolds that provide the geometric background for the applicability of our main results, Theorems 1.1-1.3. Second, in subsection 6.2, as we already noticed in the Introduction, we show that the combined assumption on the Ricci and $S$-curvatures cannot be replaced by the weighted Ricci curvature ${\bf Ric}_{N}$, in spite of the fact that the latter notion is a combination of $\mathbf{Ric}$ and $\mathbf{S}$.
@@
6.1. Funk metric spaces: applicability of Theorems 1.1-1.3
@@ A Finsler metric $F$ on $\Omega$ is called a (general) Funk metric if
Any $n$-dimensional Funk metric space $(\Omega,F,\operatorname{\mathfrak{m}}_{BH})$ endowed with the Busemann–Hausdorff measure verifies the assumptions of Theorems 1.1-1.3. In particular, the Hardy inequality, Heisenberg–Pauli–Weyl uncertainty principle and Caffarelli–Kohn–Nirenberg inequality fail on $(\Omega,F,\operatorname{\mathfrak{m}}_{BH})$.
Before giving the proof of Theorem 6.1, we recall/collect some useful properties of Funk metric spaces. We first notice that every Funk metric is projectively flat. In addition, according to Shen [32] and Li [19], a Finsler metric $F$ on $\Omega$ is a Funk metric if and only if there is a unique Minkowski norm $\phi=\phi(y)$ on $\mathbb{R}^{n}$ such that
@@
Let $(\Omega,F)$ be a Funk metric space and let $\phi=\phi(y)$ the Minkowski norm satisfying (6.1). Then the following statements hold::
@@
    $(\Omega,F)$ is forward complete but not backward complete with $\lambda_{F}(\Omega)=+\infty$. Furthermore,
@@
$\partial\Omega\ni z:=x+\frac{y}{F(x,y)}=c(s)+\frac{y}{F(c(s),y)}=x+sy+\frac{y}{F(c(s),c^{\prime}(s))},$
@@
$F(c(s),c^{\prime}(s))=\frac{F(x,y)}{1-F(x,y)s}=\frac{F(y)}{1-F(y)s},$
where $F(y)$ is $F(x,y)$ for short. Then formula (6.3) giving the distance ${d}_{F}(x_{1},x_{2})$ between $x_{1}=c(s_{1})$ to $x_{2}=c(s_{2})$ is
${d}_{F}(x_{1},x_{2})=$ $\int_{s_{1}}^{s_{2}}F(c(s),c^{\prime}(s)){\operatorname{d}}s=\int_{s_{1}}^{s_{2}}\frac{F(y)}{1-F(y)s}{\operatorname{d}}s=\ln\left[\frac{1-F(y)s_{1}}{1-F(y)s_{2}}\right]=\ln\left[\frac{\frac{1}{F(y)}-s_{1}}{\frac{1}{F(y)}-s_{2}}\right]$
@@
On the other hand, when $x_{1}=x$ and $x_{2}=\mathbf{0}$, from (6.6) we have $\lambda=\frac{1}{\phi(-x)}+1$; then (6.7) gives
@@
Let $t\mapsto c(t)$, $t\in[0,1],$ be the unit-speed geodesic from $c(0)$ to $c(1)$ (i.e., the ray $l_{c(0)c(1)}$). We extend $c$ to the point $z=c(0)+\lambda(c(1)-c(0))\in\partial\Omega$; hence, (6.1)1 implies
@@
$t=d_{F}(c(0),c(t))=d_{F}(x_{1},c(t))\geq-\ln(1-\phi(c(t)))+\ln(1-\phi(x_{1}))\rightarrow+\infty,$
as $\phi(c(t))\rightarrow 1^{-}$, which means that every unit-speed geodesic can be forward extended to $t=+\infty$. Hence, $(\Omega,F)$ is forward complete. On the other hand, (6.8) implies $d_{F}(x,\mathbf{0})\rightarrow\ln 2$ as $\phi(-x)\rightarrow 1^{-}$. Hence, ($\Omega$, $F$) cannot be backward complete because the bounded closed backward ball $\overline{B^{-}_{\mathbf{0}}(\ln 2)}=\Omega$ is not compact.
@@
with equality if and only if $\xi=\alpha\mathfrak{L}_{\phi}(x)$ for some $\alpha\geq 0$; here, $\mathfrak{L}_{\phi}$ stands for the Legendre transformation induced by $\phi$. It costs no generality to assume that $\phi^{*}(\mathfrak{L}_{\phi}(x))\leq\phi^{*}(-\mathfrak{L}_{\phi}(x))$. According to Shen [29] and Huang–Mo [11], the co-metric $F^{*}$ of $F$ is given by
@@
where $\phi^{*}$ is the co-metric of $\phi$. Hence, using (2.5), we obtain
@@
Consequently $\lambda_{F}(x)\to+\infty$ as $\phi(x)\rightarrow 1^{-}$, and hence $\lambda_{F}(\Omega)=+\infty$. Moreover,
@@
By Proposition 6.1/(iii), $(\Omega,F)$ is forward complete. Furthermore, it is established by Shen [27] that $(\Omega,F)$ is a Cartan–Hadamard manifold satisfying:
@@
Thus, in the assumptions of Theorems 1.1-1.3, we may take
@@
which ensures the validity of both (1.6) and (1.8). ∎
@@
This subsection concludes with a discussion of the interpolation metric introduced in Kristály–Rudas[18]. For $a\in[0,1]$, define $F_{a}:T\mathbb{B}^{n}\to[0,\infty)$ by
@@
This family interpolates between two fundamental metrics: the Funk metric (1.7) for $a=1$ with ${\bf K}=-1/4$ and ${\bf S}_{a}=(n+1)/2$, and the Hilbert/Klein metric for $a=0$ with ${\bf K}=-1$ and ${\bf S}_{a}=0$. Here ${\bf S}_{a}$ denotes the $S$-curvature of the Busemann–Hausdorff measure $\operatorname{\mathfrak{m}}_{a}$.
The space $(\mathbb{B}^{n},F_{a})$ is a projectively flat Finsler space of Randers-type. Its reversibility is given by $\lambda_{F_{a}}(\mathbb{B}^{n})=\frac{1+a}{1-a}$ for $a\in[0,1)$, while $\lambda_{F_{1}}(\mathbb{B}^{n})=+\infty$. Furthermore, in Kajántó–Kristály [14, Theorems 1.1 and 1.2], it is proved that for $a\in(0,1)$, up to an isometry, the flag curvature and $S$-curvature of $F_{a}$ satisfy the sharp estimates
@@
In particular, these estimates are sharp for every $a\in(0,1)$. In fact, for every $a\in[0,1)$ one has that $\inf_{(x,y)\in T\mathbb{B}^{n}}{\bf S}_{a}(x,y)=0$, and hence Theorems 1.2 and 1.3 do not apply (i.e., $0=h<k$).
Nevertheless, by combining Theorem 6.1 with elementary estimates as in [18] and the validity of famous functional inequalities on Riemannian Cartan–Hadamard manifolds [10, 16], we obtain the following sharp criterion:
@@
Let $a\in[0,1]$, and consider the interpolated Funk-type space $(\mathbb{B}^{n},F_{a},\operatorname{\mathfrak{m}}_{a})$ endowed with its natural Busemann–Hausdorff measure $\operatorname{\mathfrak{m}}_{a}$. Then the Hardy inequality, Heisenberg–Pauli–Weyl uncertainty principle, and Caffarelli–Kohn–Nirenberg inequality hold on $(\mathbb{B}^{n},F_{a},\operatorname{\mathfrak{m}}_{a})$ – with possibly not sharp constants – if and only if $a\in[0,1)$.